# ADA-Track: End-to-End Multi-Camera 3D Multi-Object Tracking with Alternating Detection and Association

### v1

@@ B.1 IDF1 Analysis
B.2 Complexity and runtime analysis
ADA-Track and TBD-Baseline require additional Edge-Augmented Cross-Attention modules compared to TBA-Baseline. We compare the number of parameters, FLOPs and runtime of ADA-Track with TBA-Baseline and TBD-Baseline in Table A2. As shown in Table A2, ADA-Track only adds about $6.7\%$ parameters, $2.9\%$ FLOPs, $4.1\%$ inference time to TBA, and even less compared to TBD. Therefore, the additional computational overhead of ADA-Track is minimal. Despite this slight increase in complexity, the performance gain of ADA-Track is much more significant, e.g. $17.6\%$ higher AMOTA than TBA-Baseline with DETR3D (see Table 1).
B.3 Ablation Studies
@@
Center 0.366 1.392 0.510 0.325 916
None 0.341 1.442 0.474 0.300 1442
Appearance 0.358 1.417 0.510 0.318 1088
Box 0.378 1.391 0.507 0.343 981
Table A3: Ablation study on appearance and geometric features to build the relative positional encoding in the edge-augmented cross-attention. Our choice is Box (last row) which uses all the box parameters to build geometric-based relative position encoding. Center denotes that only box centers are used. None denotes no relative position encoding. Appearance uses the query feature differences as appearance-based relative position encoding.
@@
0.346 1.421 0.498 0.310 1275
0.378 1.391 0.507 0.343 981
Table A4: Ablation study on the iterative refinement of edge features over decoder layers.
@@
0.367 1.404 0.504 0.320 936
0.364 1.411 0.503 0.317 899
0.378 1.388 0.513 0.344 904
0.378 1.391 0.507 0.343 981
Table A5: Ablation study on computing attention across different query types in self-attention.
Appearance and geometry cues for association
We investigate the role of appearance and geometric cues in the learnable association module based on edge-augmented cross-attention. As shown in Table A3, using only the center position (first row) instead of the complete box parameters (fourth row) leads to 1.2%P decrease in AMOTA, which underscores the significance of leveraging the entire box information for robust data association. If geometric features are excluded (second row), the zero-initialized edge features are refined exclusively through appearance-based query features layer-by-layer, resulting in a substantial performance drop across all metrics when compared to the use of geometric-based edge features. This observation highlights the usage of geometric features in enhancing the model’s ability to distinguish between object instances. Even without geometric features, using relative positional encodings derived from query feature differences (third row) yields a notable AMOTA increase of 1.7%P compared to scenarios without relative positional encoding (second row). This finding shows the importance of the exact edge feature encoding in the model architecture of edge-augmented cross-attention.
Edge feature refinement
Table A4 illustrates the effectiveness of the iterative refinement of edge features over decoder layers. In the case where edge features remain independent within each decoder layer (first row), there is a notable decrease in AMOTA by 3.2%P when compared to scenarios where edge feature refinement occurs across layers (second row). This experiment shows the potential for iterative optimization of data association across decoder layers, aligning with the fundamental design of our architecture. In addition, since the edge features also participate in the query feature update in the edge-augmented cross-attention, the refinement of the edge features itself also contributes to the iterative refinement of query representations, which also improves the overall performance.
Masked self-attention
The self-attention layer in ADA-Track facilitates temporal modeling between queries. As shown in Table A5, when we use only the attention from track to detection queries in the self-attention layer (third row), we observe no changes in AMOTA compared to our default setting (fourth row). The other metrics are even slightly better, indicating that this setting is slightly preferable compared to the default setting that we used in all other experiments. Conversely, using only the attention from detection to track queries (second row) results in a noticeable drop of 1.4%P in AMOTA. The results are similar when attention is not computed across different query types (first row). These results show that self-attention is important for enhancing detection queries, enabling detection queries to incorporate information from past tracks and frames.
Association module
Table A6 compares Edge-Augmented Cross-Attention with alternative association modules. We replace the Edge-Augmented Cross-Attention with association networks utilizing the difference or concatenation of detection and track query features (node features). In both cases, we use an MLP and sigmoid to obtain the association scores $S$ as before. Using only the difference or concatenation of node features results in a significant performance drop, highlighting the necessity of using explicit edge features and the effectiveness of Edge-Augmented Cross-Attention in data association.
Robustness against appearance change
One of the biggest challenges of MOT in autonomous driving is ego-motion, where the camera mounting points move with the ego-vehicle, leading to appearance changes of observed objects due to varying observation angles. We evaluate the sequences where the average speed of the ego-vehicle is $\geq 5m/s$, indicating significant appearance changes in observed objects due to ego-motion. As shown in Table A7, ADA-Track achieves similar AMOTA for sequences with ego-motion ($\geq 5m/s$) compared to all scenes ($\geq 0m/s$). Some secondary metrics are even better for the sequences with ego-motion.
@@
Node Difference 0.355 1.421 0.483 0.314 1249
Node Concatenation 0.349 1.421 0.451 0.306 977
Edge-Aug. Cross Attn. 0.378 1.391 0.507 0.343 981
Table A6: Ablation study on different association modules.
@@
$\geq 5m/s$ 0.377 1.378 0.512 0.355 589
$\geq 0m/s$ 0.378 1.391 0.507 0.343 981
Table A7: Abaltion study on the impact of ego-motion.
Number of queries
Table A8 shows the comparison of varying the number of detection queries $N_{D}$, where AMOTA increases until $N_{D}=300$. Using fewer detection queries typically causes missed detections and lower detection performance, inevitably affecting the tracking performance. However, continuing to increase the number of detection queries does not yield further improvements. This is attributed to the risk of introducing an imbalance in the classification for the data association task, potentially reducing the association performance. As a result, we opt for $N_{D}=300$ as the default setting.
@@
100 0.341 1.437 0.461 0.301 1213
200 0.370 1.397 0.471 0.327 847
300 0.378 1.391 0.507 0.343 981
400 0.377 1.395 0.486 0.329 805
500 0.377 1.397 0.514 0.335 919
Table A8: Ablation study on the number of detection queries $N_{D}$.
@@
2 0.371 1.408 0.507 0.332 1000
5 0.378 1.386 0.518 0.331 922
10 0.378 1.391 0.507 0.343 981
20 0.377 1.389 0.504 0.336 920
50 0.365 1.399 0.516 0.327 911
Table A9: Ablation study on the association loss weight $\lambda_{\text{asso}}$.
@@
0.0 0.358 1.404 0.496 0.323 1659
0.5 0.359 1.409 0.506 0.323 1118
1.0 0.378 1.391 0.507 0.343 981
1.5 0.363 1.384 0.516 0.326 984
2.0 0.366 1.406 0.465 0.322 952
Table A10: Ablation study on the focusing parameter $\gamma$ of the association loss.
Association loss weight
We evaluate the weight of the association loss $\lambda_{\text{asso}}$ in Table A9. Using $\lambda_{\text{asso}}=5$ and $\lambda_{\text{asso}}=10$ yield the same AMOTA of 0.378. Lower or higher association loss weights result in performance drops with different ranges, which can be attributed to the imbalance of the multi-task training. We choose $\lambda_{\text{asso}}=10$ as default.
Focusing parameter in association loss
We use the focal loss as the association loss $\mathcal{L}_{\text{asso}}$ with a focusing parameter of $\gamma=1.0$. This choice is validated in Table A9, where lowering or raising the focusing parameter $\gamma$ results in a notable decrease of AMOTA, ranging from 1.2%P to 2.0%P. Therefore $\gamma=1.0$ is a reasonable choice for effectively controlling the class imbalance in the data association task.
@@
2 0.361 1.421 0.495 0.337 1246
3 0.369 1.404 0.496 0.338 1070
4 0.375 1.395 0.503 0.340 1072
5 0.378 1.391 0.507 0.343 981
6 0.373 1.392 0.501 0.337 934
Table A11: Ablation study on the duration of keeping unmatched tracks $T_{d}$.
@@
0.0 0.203 1.421 0.347 0.213 1575
0.1 0.248 1.409 0.380 0.246 1827
0.2 0.315 1.386 0.462 0.280 1520
0.3 0.365 1.379 0.495 0.327 1217
0.4 0.378 1.391 0.507 0.343 981
0.5 0.364 1.424 0.489 0.337 869
Table A12: Ablation study on the score threshold for track spawning $\tau_{\text{new}}$.
Hyperparameters during inference
During inference, we use two hyperparameters: the number of frames until unmatched tracks are kept and the score threshold $\tau_{\text{new}}$ to spawn new tracks. We evaluate both hyperparameters in Table A11 and Table A12. As shown in Table A11, low values of $T_{d}$ cause a significant performance drop caused by the insufficient handling of occluded objects. The AMOTA peaks at $T_{d}=5$ and a higher value again leads to a decrease of AMOTA, which might keep too many tracks and cause a higher class imbalance in the data association. As for the score threshold $\tau_{\text{new}}$, when setting $\tau_{\text{new}}\leq 0.3$, the tracker initializes excessive noisy detections, which results in a significant performance drop. We choose $\tau_{\text{new}}=0.4$ as the default value.

### v2

@@ 5.5.3 Track update
Box update Both track and detection queries predict bounding boxes, resulting in two candidates for each instance when two queries are associated. We use the box from the detection side to represent this associated pair and validate this choice in Table VII. As shown in the first row (Exp. A), selecting the track box results in a decrease of 0.4%P AMOTA and $1.4\%$P recall compared to our approach (Exp. C), which shows the better quality of detection boxes. One could argue that detection queries are further refined in the edge-augmented cross-attention but track queries are not. In the second row of Table VII, we additionally reverse the edge-augmented cross-attention by using tracks as queries and detections as keys. This leads to an AMOTA decrease of 1.3%P and a substantial recall decrease of $4.1\%$P. This finding also verifies the necessity of the decoupled queries for decoupled tasks, where using detection queries to locate objects and associate them to track queries is more effective than relying on track queries to locate objects.
@@
Center 0.388 1.365 0.504 0.344 895
None 0.378 1.388 0.523 0.340 1123
Appearance 0.376 1.357 0.515 0.342 824
Box 0.391 1.364 0.524 0.343 842
Table VIII: Ablation study on choices of the relative positional encoding in the edge-augmented cross-attention. Our choice is Box (last row) which uses all the box parameters to build geometric-based relative position encoding. Center denotes that only box centers are used. None denotes no relative position encoding. Appearance uses the query feature differences as appearance-based relative position encoding.
@@
0.376 1.376 0.492 0.334 954
0.391 1.364 0.524 0.343 842
Table IX: Ablation study on the iterative refinement of edge features over decoder layers.
5.5.4 Edge features
Appearance and geometry cues for association We investigate the role of appearance and geometric cues in the learnable association module based on edge-augmented cross-attention. As shown in Table VIII, using only the center position (first row) instead of the complete box parameters (fourth row) leads to a slight decrease in AMOTA, which underscores the significance of leveraging the entire box information for robust data association. If geometric features are excluded (second row), the zero-initialized edge features are refined exclusively through appearance-based query features layer-by-layer, resulting in a substantial performance drop across all metrics when compared to the use of geometric-based edge features. Still without geometric features, using relative positional encodings derived from query feature differences (third row) yields a notable AMOTA decrease of 1.5%P compared to the default setting (last row). This observation highlights the usage of geometric features in enhancing the model’s ability to distinguish between object instances.
Edge feature refinement With an iterative edge feature refinement, we observe a notable decrease in AMOTA by 1.5%P compared to independent edge features within each decoder layer in Table IX. This experiment shows the potential for iterative optimization of data association across decoder layers, aligning with the fundamental design of our architecture. In addition, since the edge features also participate in the query feature update in the edge-augmented cross-attention, the refinement of the edge features itself also contributes to the iterative refinement of query representations, which again confirms our architecture design.
[…]
Node Difference 0.366 1.386 0.500 0.327 1201
Node Concatenation 0.364 1.388 0.500 0.327 1143
Edge-Aug. Cross Attn. 0.391 1.364 0.524 0.343 842
Table X: Ablation study on different association modules.
@@ 5.5.4 Edge features
0.381 1.364 0.511 0.341 856
0.386 1.358 0.502 0.349 805
0.388 1.366 0.526 0.347 968
0.391 1.364 0.524 0.343 842
Table XI: Ablation study on computing attention across different query types in self-attention.
@@
2 0.378 1.369 0.510 0.346 902
5 0.387 1.360 0.513 0.350 829
10 0.391 1.364 0.524 0.343 842
20 0.388 1.360 0.531 0.351 853
50 0.384 1.364 0.522 0.345 875
Table XII: Ablation study on the association loss weight $\lambda_{\text{asso}}$.
5.5.5 Learned association module
Table X compares edge-augmented cross-attention with alternative node-only data association modules in ADA-Track. We replace the edge-augmented cross-attention with association networks utilizing the difference or concatenation of detection and track query features (node features). In both cases, we use an MLP and sigmoid to obtain the association scores $S$ as before. Using only the difference or concatenation of node features results in a significant performance drop, highlighting the necessity of using explicit edge features and the effectiveness of edge-augmented cross-attention in data association.
5.5.6 Masked self-attention
The self-attention layer in ADA-Track facilitates temporal modeling between queries. As shown in Table XI, without interaction between track and detection queries (first row), the model achieves only an AMOTA of 0.381. In contrast, allowing attention in a single direction, either from detection to track queries or vice versa, improves overall performance. Notably, enabling attention from track to detection queries has a more significant positive impact. When attention is applied in both directions, the AMOTA improves further, underscoring the importance of bidirectional attention in enhancing query representations and overall performance.
5.5.7 Association loss weight
We evaluate the weight of the association loss $\lambda_{\text{asso}}$ in ADA-Track in Table XII. Using $\lambda_{\text{asso}}=10$ yields the highest AMOTA of 0.391. Lower or higher association loss weights result in performance drops, which can be attributed to the imbalance of the multi-task training. Therefore, we choose $\lambda_{\text{asso}}=10$ as default.
5.6 Ablation Study on ADA-Track++
We further provide more ablation studies on the improved version ADA-Track++ and compare it with ADA-Track.
@@ 5.6 Ablation Study on ADA-Track++
ADA- Track 100 0.371 1.393 0.505 0.333 1062
200 0.379 1.380 0.504 0.340 837
300 0.391 1.364 0.524 0.343 842
400 0.389 1.365 0.524 0.347 889
500 0.389 1.363 0.529 0.351 779
ADA- Track++ 100 0.372 1.381 0.508 0.330 829
200 0.388 1.351 0.515 0.343 844
300 0.398 1.336 0.544 0.357 774
400 0.400 1.344 0.538 0.360 777
500 0.397 1.334 0.526 0.359 734
Table XIII: Impact of the number of detection queries $N_{D}$ on ADA-Track and ADA-Track++.
@@
ADA- Track 0.01 0.368 1.379 0.495 0.312 1286
0.1 0.385 1.365 0.523 0.336 910
0.2 0.388 1.358 0.496 0.335 760
0.3 0.391 1.364 0.524 0.343 842
0.4 0.382 1.365 0.512 0.339 1029
ADA- Track++ 0.01 0.388 1.338 0.525 0.334 808
0.1 0.394 1.339 0.542 0.351 716
0.2 0.397 1.337 0.530 0.354 689
0.3 0.398 1.336 0.544 0.357 774
0.4 0.398 1.338 0.531 0.369 859
Table XIV: Impact of the affinity threshold $\tau_{S}$ on ADA-Track and ADA-Track++.
@@
ADA- Track 2 0.378 1.389 0.514 0.343 1156
3 0.386 1.370 0.523 0.342 952
4 0.386 1.369 0.516 0.343 873
5 0.391 1.364 0.524 0.343 842
6 0.392 1.355 0.517 0.343 803
ADA- Track++ 2 0.391 1.358 0.533 0.355 1074
3 0.397 1.339 0.536 0.355 886
4 0.399 1.336 0.517 0.358 791
5 0.398 1.336 0.544 0.357 774
6 0.398 1.337 0.540 0.357 741
Table XV: Impact of the maximum track age $T_{D}$ on ADA-Track and ADA-Track++.
5.6.1 Number of queries
We evaluate the impact of varying the number of detection queries $N_{D}$ on both ADA-Track and ADA-Track++, as shown in Table XIII. Reducing the number of detection queries typically leads to missed detections and a decline in detection performance, which subsequently affects the tracking performance in both variants. For ADA-Track, we observe that AMOTA increases with the number of detection queries up to $N_{D}=300$. However, further increasing $N_{D}$ does not result in additional improvements. In contrast, ADA-Track++ reaches its peak AMOTA at $N_{D}=400$, indicating its improved ability to accurately associate tracks with a larger number of detections. This result underscores the effectiveness of the auxiliary token introduced in ADA-Track++, which reduces the risk of conflicting affinity scores as discussed in Section 4. Notably, ADA-Track++ consistently exhibits lower IDS across all $N_{D}$ values compared to ADA-Track, indicating its enhanced ability to reduce confusion in data association.
5.6.2 Affinity threshold
During inference, our method filters the estimated affinity score $S$ using a threshold $\tau_{S}$ before applying the Hungarian Matching (HM). Table XIV presents the impact of this threshold on both ADA-Track and ADA-Track++. For both variants, the optimal AMOTA is achieved with $\tau_{S}=0.3$, while any increase or decrease in the threshold leads to a more significant decline in performance of ADA-Track than ADA-Track++. For instance, at the threshold $\tau_{S}=0.01$, where the AMOTA of ADA-Track drops remarkably to 0.368, ADA-Track++ maintains a higher AMOTA of 0.388 and an IDS of 808. This demonstrates that the affinity scores $S$ in ADA-Track++ are less noisy, largely due to the introduction of the auxiliary token, which effectively mitigates confusion in the attention scores and enhances the robustness.
5.6.3 Maximum track age
@@
sigmoid 0.390 1.345 0.532 0.350 715
softmax 0.360 1.346 0.503 0.314 1042
softmax 0.383 1.338 0.528 0.332 1180
sigmoid 0.398 1.336 0.544 0.357 774
Table XVI: Ablation study on the association loss and the activation function of the affinity scores $S$ for ADA-Track++.
@@
0.02 0.388 1.344 0.533 0.351 688
0.05 0.387 1.349 0.526 0.346 748
0.1 0.398 1.336 0.544 0.357 774
0.2 0.394 1.351 0.528 0.349 677
0.5 0.385 1.345 0.533 0.346 682
Table XVII: Ablation study on the categorical cross entropy loss weight $\lambda_{\text{CE}}$ in $\mathcal{L}_{\text{asso}}$.
[…]
w/o self-attn 0.395 1.340 0.517 0.349 695
w/o prop 0.396 1.356 0.539 0.358 786
with prop 0.398 1.336 0.544 0.357 774
Table XVIII: Ablation study on the auxiliary token in ADA-Track++. Our choice (last row) is to propagate the auxiliary token over time stamps and let it participate in the query-to-query self-attention. w/o self-attn denotes that the auxiliary token is not involved in the self-attention. w/o prop denotes that the auxiliary token is initialized every frame.
5.6.4 Loss function and activation
[…]
Table XVII evaluates the impact of the categorical cross entropy loss weight $\lambda_{\text{CE}}$ of ADA-Track++, where $\lambda_{\text{CE}}=0.1$ leads to the best overall performance, especially AMOTA and Recall. Therefore we pick $\lambda_{\text{CE}}=0.1$ as the default value.
5.6.5 Auxiliary token feature
In Table XVIII, we evaluate different choices for the feature representation of the auxiliary token. Compared to our default setting (last row), initializing the auxiliary token at each frame (second row) results in a slight AMOTA decrease to 0.396. Similarly, when the auxiliary token is excluded from the query-to-query self-attention, a slight decrease in AMOTA is also observed. These results confirm our design choice, which allows the auxiliary token to interact more extensively with other queries to capture more scene-level information, leading to a more robust feature representation and improved overall performance.
6 Conclusion
We presented a novel query-based multi-camera 3D multi-object tracking approach. We observed that decoupling the detection and association tasks while simultaneously leveraging the synergies between these tasks is the key to accomplish high-quality tracking. In line with this finding, we proposed a paradigm that conducts both, detection and association, in an alternating manner. In addition, we proposed a learned association module based on edge-augmented cross-attention which can be seamlessly integrated into any query-based decoder. To further improve the association process, we incorporated an auxiliary token as an additional association target, reducing the risk of confusion caused by softmax attention normalization. Extensive experiments show the effectiveness of our approach compared to other paradigms, while achieving promising performance on the nuScenes tracking benchmark. Moreover, our method is compatible with 3D detectors and can easily incorporate insights from other tracking methods, making it highly flexible and adaptable to different setups. We hope that our method can inspire future research in the community and serve as an established baseline for 3D multi-object tracking.
Acknowledgments